OpenStudy (anonymous):

Find \[\lim_{x \rightarrow \pi} \cot(x)\] determine whether it is a positive or negative limit.

OpenStudy (astrophysics):

Hint: \[\cot(x) = \frac{ 1 }{ \tan(x) }\]

OpenStudy (anonymous):

so it'd be a positive limit correct?

OpenStudy (empty):

The limit doesn't exist, because if you approach from the left it will be negative and if you approach from the right it will be positive.

OpenStudy (astrophysics):

The limit actually doesn't exist

OpenStudy (astrophysics):

Yeah haha

OpenStudy (astrophysics):

from - you get - infinity and + you get infinity

OpenStudy (empty):

Look at this graph as you approach \(\pi\) http://www.biology.arizona.edu/biomath/tutorials/trigonometric/graphics/trig_cotan.gif

OpenStudy (astrophysics):

\[\lim_{x \rightarrow \pi^+} cotx = - \infty ~~~~~~ \lim_{x \rightarrow \pi ^-} cotx = \infty \]

OpenStudy (anonymous):

what if it were cot(2x) instead?

OpenStudy (astrophysics):

You should try them out yourself